Convert to augmented matrix.
x+y+z=1
-x=2+z-3y
z=x+y

Answer by tommyt3rd(5050)

x+y+z=1
-x=2+z-3y
z=x+y

x+y+z=1
-2=x+z-3y
0=x+y-z



with a vertical bar between columns 3 and 4
:)